OpenStudy (abbycross167):

Can someone please show me the steps to answer this question? A line with slope 1/3 passes through the point (3, 6). Which point also lies on this line? a. (6, 3) b. (7, 6) c. (-3, -3) d. (-6, 3)

OpenStudy (anonymous):

y=slope * x coordinate + y coordinate so y= 1/3(3) + 6 simplify: y=1+6 so y=7. Without having to solve for y you know it's (7,6) because it's the only option with a 7